OpenStudy (dariyana):

A ladder leans against a house at a 60° angle to the ground. If the ladder extends to a length of 132 inches, what is the height of the house rounded to the nearest hundredth of an inch? Type the numeric answer only in the box below.

OpenStudy (lgbasallote):

 does that help?

OpenStudy (dariyana):

136.83 rite

OpenStudy (lgbasallote):

google calculator says no...

OpenStudy (anonymous):

\[\Large \sin(60)=\frac{perpendicular}{hyp}\]

OpenStudy (anonymous):

@dariyana \[\sin 60 = x/132\] therefore x= 132 *sin60

OpenStudy (lgbasallote):

check to see that your calculator is in degrees not radians

OpenStudy (lgbasallote):

you should get something lower than that...132 is supposed to be the longest side

OpenStudy (lgbasallote):

so you know x should be less than 132

OpenStudy (dariyana):

114.32

OpenStudy (lgbasallote):

right